OpenStudy (anonymous):

Find a formula for the general term an of the sequence (starting with a1): {-4/25,5/36,-6/49,7/64,...}

OpenStudy (anonymous):

\[\begin{align*} a_1&=\frac{(-1)^1(1+3)}{(1+4)^2}=-\frac{4}{25}\\ a_2&=\frac{(-1)^2(2+3)}{(2+4)^2}=\frac{5}{36}\\ \vdots\\ a_n&=\cdots \end{align*}\]